OpenStudy (andrewkaiser333):

Please help me!!! The factors of 5y2 − 2y − 7 are (5y − 7) and ( ).

jhonyy9 (jhonyy9):

just you need dividing this quadratic by (5y-7) and will get easy the right second factor

OpenStudy (andrewkaiser333):

how would i do that if idk y?

jhonyy9 (jhonyy9):

5y^2 -2y -7 / 5y -7 = y +1 5y^2 -7y --------- 0 5y -7 5y -7 -------- 0 0 hope helped

OpenStudy (will.h):

5y^2 -2y -7 The 2 numbers when multipleD = -35 and when added = -2 is -7 and 5 So re express (5y^2 + 5y)( -7y -7) GCF of the 1st bi is 5y 5y(y + 1) GCF of the 2nd bi is -7 -7(y + 1) The final answer is (5y - 7)( y + 1)

jhonyy9 (jhonyy9):

hi Will ! Saluti o Ciao !

OpenStudy (will.h):

Hey Andrew :) I liked your way tho

OpenStudy (andrewkaiser333):

thnks guys!

jhonyy9 (jhonyy9):

ok. good luck and yw.